1 pts Question 24 The current yield on a $6,000, 10 percent coupon bond selling for $5,000 is 5% 10% 12% 15%
Answer #1

Annual coupon=6000*10%=$600

Hence current yield=Annual coupon/Current price

=600/5000

which is equal to

=12%